# Prompting Depth Anything for 4K Resolution Accurate Metric Depth Estimation

### v2

@@ 3.4 Implementation Details
We utilize the ViT-large model as our backbone model. The shallow convolutional network comprises two convolutional layers with a kernel size of 3 and a stride of 1. More details can be found in the supp. Detailed running time analysis can be found in Sec. 4.3.
@@
The irregular range of input depth data can hinder network convergence. To address this, we normalize the LiDAR data using linear scaling to the range [0, 1], based on its minimum and maximum values. The network output is also normalized with the same scaling factor from LiDAR data, ensuring consistent scales and facilitating easier convergence during training.
@@
Table 1: Quantitative comparisons on ARKitScenes dataset. The terms Net., Post. and w/o LiDAR refer to the LiDAR depth usage of models, where “Net.” denotes network fusion, “Post.” indicates post-alignment using RANSAC, and “w/o LiDAR” means the output is metric depth. Methods marked with are finetuned with their released models and code on ARKitScenes [3] and ScanNet++ [78] datasets.
L1 $\downarrow$ RMSE $\downarrow$ AbsRel $\downarrow$ $\delta_{0.5}$uparrow Acc $\downarrow$ Comp $\downarrow$ Prec $\uparrow$ Recall $\uparrow$ F-score $\uparrow$
MSPF 0.0326 0.0975 0.0226 0.9674 0.0772 0.0695 0.6738 0.7761 0.7133
@@
D.P. 0.0353 0.0983 0.0242 0.9657 0.0820 0.0747 0.6431 0.7234 0.6734
Table 2: Quantitative comparisons on ScanNet++ dataset. The terms Net., Post. and w/o LiDAR refer to the LiDAR depth usage of models as the last table. Methods marked with are finetuned with their released code on ARKitScenes [3] and ScanNet++ [78] datasets.
@@
Table 3: Quantitative ablations on ARKitScenes and ScanNet++ datasets. Please refer to Sec. 4.3 for detailed descriptions.
4 Experiments
4.1 Experimental Setup
We mainly conduct experiments on the HyperSim synthetic dataset [51] and two real-world datasets: ScanNet++ [78] and ARKitScenes [3], which provide iPhone RGB-LiDAR data ($192\times 256$ resolution) and annotated depth from a high-power LiDAR ($1440\times 1920$ resolution). We follow the suggested training and evaluation protocol in [3] for ARKitScenes, where 40K images are used for training and 5K images for evaluation. For the ScanNet++ dataset, we randomly select 20 scenes from its 50 validation scenes, amounting to approximately 5K images for our validation and the training set are from its 230 training scenes, containing about 60K images. To ensure a fair comparison, we additionally train a model with HyperSim training set to achieve zero-shot testing on ScanNet++ and ARKitScenes datasets. Besides depth accuracy metrics, we also report the TSDF reconstruction results of our method on ScanNet++, which reflects the depth consistency. We describe the details of the evaluation metrics in the supp.
4.2 Comparisons with the State of the Art
[…]
4.3 Ablations and Analysis
@@
We assess its importance with two experiments: 1) Removing the prompting. Tab. 3(b) shows a significant performance drop. 2) Removing the foundation model initialization [76]. Tab. 3(c) shows a noticeable performance decline.
@@
We study different designs: AdaLN, Cross-attention, and ControlNet as discussed in Sec. 3.2. Tab. 3(d,e,f) reveals that ControlNet performs best but still falls short of our method.
@@
We initially incorporate ARKitScenes data, which only enhances performance on ARKitScenes (Tab. 3(g)). Then we add ScanNet++, which improves results on both ARKitScenes and ScanNet++ (Tab. 3(h)). However, the depth visualization remains less than ideal (Fig. 3). Tab. 3(i) show that direct supervision with pseudo GT depth from reconstruction methods decreases performance. Ultimately, employing the edge-aware depth loss that utilizes pseudo GT depth and FARO annotated GT achieves comparable performance with  Tab. 3(h) but with superior thin structure depth performance as shown in Fig. 3. We provide more qualitative ablation results in the supp.
Figure 6: Outdoor reconstruction by taking the vehicle LiDAR as metric prompt. Please refer to the supp. for more video results.
@@
Our model with ViT-L runs at 20.4 FPS for an image resolution of $768\times 1024$ on a A100 GPU. As ARKit6 supports 4K image recording, we test our model at a resolution of $2160\times 3840$ and achieve 2.0 FPS. Note that our model can also be implemented with ViT-S, where the corresponding speeds are 80.0 and 10.3 FPS. More testing results can be found in the supp.
Figure 7: Zero-shot testing on diverse scenes.
4.4 Zero-shot Testing on Diverse Scenes
Although our model is trained on indoor scenes, it generalizes well to various scenarios, including new rooms, gyms with thin structures, poorly lit museums, human and outdoor environments, as shown in Fig. 7, highlighting the effectiveness of prompting a depth foundation model. Please refer to the supp. for video results.
4.5 Application: 3D Reconstruction
Our consistent and scale-accurate depth estimation benefits the indoor 3D reconstruction as shown in Tabs. 2 and 5. Besides, the prompt of our model can be easily replaced with vehicle LiDAR, which enables our model to achieve large-scale outdoor scene reconstruction as shown in Fig. 6. We detail the setup and include more video results for dynamic streets in the supp.
4.6 Application: Generalized Robotic Grasping
@@
We set up a robotic platform to test our model in generalized robotic manipulation (Fig. 8), which typically requires depth or RGB as observations. Good depth estimation enhances the generalization ability because it accurately describes the 3D information of surroundings [82, 27]. Specifically, we train an ACT policy [85] to grasp various objects into the box, using different types of input signals such as RGB, LiDAR, and depth data from our model. We empirically find that our model generalizes well to unseen objects like transparent and specular objects when trained on diffusive objects, outperforming RGB and LiDAR inputs as shown in Tab. 4. This is because RGB is dominated by color, which leads to poor generalization across objects, and the iPhone LiDAR depth is noisy and lacks the capability to perceive transparent objects. Please refer to the supp. for detailed setup descriptions and videos.
@@
Table 4: Grasping success rate on various objects. Three numbers indicate objects placed at near, middle, and far positions. The grasping policy is trained on diffusive and tested on all objects.
5 Conclusion and Discussions
@@
Appendix A Additional Discussions
Figure 10: Generalizability to different resolutions. Our model can infer depth for images of different resolutions from 512p to 2160p.
A.1 Generalizability to Different Resolutions
This section discusses the generalization capability of our model across different image and lidar depth resolutions provided by ARKit4 and ARKit6. ARKit4 captures images at a maximum resolution of $1440\times 1920$ at 60Hz and lidar depth at $192\times 256$, while ARKit6 captures images at a maximum resolution of $3024\times 4032$ at 30Hz and lidar depth at $240\times 320$. Both ScanNet++ [78] and ARKitScenes [3] are collected using ARKit4. Although our model is trained using ScanNet++ and ARKitScenes data at a resolution of $1440\times 1920$, we find that it generalizes well to ARKit6 images and depth at different resolutions. As shown in Fig. 10, we include a comparison of depth estimation for images of different resolutions, with an image resolution of $2160\times 3840$ and a lidar depth resolution of $144\times 256$, captured from the ARKit6 API.
A.2 Why Do We Need Synthetic Data?
The advantages of synthetic data include high-quality ground truth depth, which has been crucial for the success of many recent depth estimation works [76, 30, 19, 8]. We also utilize synthetic data to achieve high-quality depth estimation results. Furthermore, the availability of real data with lost-cost LiDAR and high-power LiDAR is currently limited [78, 3], primarily to indoor scenes, while synthetic data can further enhance diversity; for instance, our experiments have shown that including human synthetic data [81] improves our method’s generalization to human subjects.
Figure 11: Effects of using real data.
A.3 Why Do We Need Real Data?
Training with real data can further address the inability of synthetic LiDAR simulation to replicate LiDAR noise patterns, thereby enhancing depth estimation capabilities. By utilizing synthetic data, we have achieved preliminary results. However, as demonstrated by the quantitative experiments in the main paper, the use of real data further enhances the performance. Here, we include additional qualitative results in Fig. 11, which show that real data is beneficial because LiDAR simulation methods cannot fully replicate the noise of real LiDAR.
  ARKitScenes ScanNet++
  (a) Depth Any. as foundation 0.0132 0.0115 0.0699 0.0616 0.7619
(b) Depth Pro as foundation 0.0169 0.0150 0.0754 0.0676 0.7202
(c) Depth Pro 0.1225 0.1038 0.0904 0.0760 0.6187
Table 5: Additional quantitative ablations. Please refer to Sec. A.4 for detailed descriptions.
A.4 Replacing Depth Foundation Models
Since our model is a general design for DPT, it can be easily adapted to other depth foundation models that also utilize the DPT structure, such as Depth Pro [8]. Our experiments demonstrate that it significantly enhances the performance of Depth Pro, as shown in Tab. 5-(b,c), although it does not outperform our choice of Depth Anything Tab. 5-(a).
Appendix B Additional Results
Figure 12: Visualization results of simulated LiDAR. “Interp. Simu.” is the proposed interpolation method, which is interpolated from sparse anchors depth. This method effectively simulates the noise of real LiDAR data. We also provide the naive downsampled simulated LiDAR for comparison.
Visualization results of simulated LiDAR.
We provide the visualization results of our simulated LiDAR in Fig. 12.
Figure 13: ZipNeRF depth of different training frames. Training with resampled frames removing blurred frames leads to a better ZipNeRF reconstruction.
ZipNeRF reconstruction results.
High-quality and dense observations are essential for effective 3D reconstruction. However, the iPhone data from ScanNet++ [78] frequently exhibits motion blur. To address this, we resample videos from ScanNet++ to remove blurring frames. Specifically, we calculate the variance of Laplacians for each image to assess its sharpness and use the sharpness score to select frames. For a 60fps video, we select one frame every 30 frames, ensuring no repeated selection within any 6 consecutive frames, and guarantee at least one selection within every 2 seconds. We find that this method significantly reduces motion blur and leads to a better ZipNeRF reconstruction as shown in Fig. 13. Additionally, we utilize both the DSLR and iPhone data released by ScanNet++ to optimize ZipNeRF, which substantially improved our experimental results. Training ZipNeRF on ScanNet++ data required approximately $280\times 2.5\times 8$ GPU hours. We will release our processed data to benefit the research community.
Figure 14: Illustration of different depth annotation types. Please refer to Appendix B for more descriptions.
Illustration of different annotation types.
We provide an illustration of different annotation types in Fig. 14. Here we clearly observe the issues and advantages of different depth annotation types. The GT depth in ScanNet++ is annotated using a FARO scanned mesh. Due to the presence of many occlusions in the scene, the scanned mesh is incomplete, resulting in depth maps with numerous holes and poor edge quality. The pseudo GT depth annotated using NeRF reconstruction has accurate edges but performs poorly in planar regions. Therefore, an edge-aware loss is proposed to merge their advantages.
Appendix C More Details
C.1 Details about Our Model
We employ the ViT-large model from Depth Anything v2 [76] as our backbone model. The shallow convolutional network consists of two convolutional layers, each with a kernel size of 3 and a stride of 1, utilizing ReLU as the non-linear activation function. The zero-initialized projection layer is a $1\times 1$ convolutional layer. For training on the ScanNet++ [78] dataset, we apply the loss function proposed in the main paper. For training on the ARKitScenes [3] dataset, we exclusively use the L1 loss. For training on synthetic [51] data, we employ both gradient and pixel-wise L1 loss simply from ground-truth depth supervision.
Figure 15: Illustrations of our method and optional designs. Please refer to Sec. C.2 for more details.
C.2 Optional Design Details
As mentioned in the main paper, in addition to the proposed design, we also explore optional designs including AdaLN [45], Cross-attention [60], and ControlNet [83]. We include a figure to illustrate these designs in Fig. 15. Our experiments (Tab. 3 in the main paper) show that ControlNet performs the best among these alternatives, but it is still not as effective as our proposed design. The plausible reason is that they are designed to integrate cross-modal information (e.g., text prompts), which does not effectively utilize the pixel alignment characteristics between the input low-resolution depth and the output depth. We also combine the proposed design with ControlNet to investigate potential further improvements. However, no additional improvements are observed (ours vs. combination are 0.730 vs. 0.731 in terms of F-score metric on ScanNet++), but the computational costs increase. Therefore, we keep the proposed design in the final version.
C.3 Evaluation Metrics
For depth metrics, we report L1, RMSE, AbsRel and $\delta_{0.5}$. Their definitions can be found in Tab. 6.
Metric Definition
L1 $\frac{1}{N}\sum_{i=1}^{N}|\mathbf{D}_{i}-\hat{\mathbf{D}}_{i}|$
RMSE $\sqrt{\frac{1}{N}\sum_{i=1}^{N}(\mathbf{D}_{i}-\hat{\mathbf{D}}_{i})^{2}}$
AbsRel $\frac{1}{N}\sum_{i=1}^{N}|\mathbf{D}_{i}-\hat{\mathbf{D}}_{i}|/\mathbf{D}_{i}$
$\delta_{0.5}$ $\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}\left(\max\left(\frac{\mathbf{D}_{i}}{\hat{\mathbf{D}}_{i}},\frac{\hat{\mathbf{D}}_{i}}{\mathbf{D}_{i}}\right)<1.25^{0.5}\right)$
Table 6: Depth metric definitions. $\mathbf{D}$ and $\hat{\mathbf{D}}$ are the ground-truth and predicted depth, respectively. $\mathbb{I}$ is the indicator function.
For reconstruction metrics, we report Acc, Comp, Prec, Recall, F-score. Their definitions can be found in Tab. 7. We use a voxel size of 0.04m for TSDF reconstruction.
Metric Definition
Acc $\mbox{mean}_{p\in P}(\min_{p^{*}\in P^{*}}||p-p^{*}||)$
Comp $\mbox{mean}_{p^{*}\in P^{*}}(\min_{p\in P}||p-p^{*}||)$
[…]
F-score $\frac{2\times\text{Perc}\times\text{Recal}}{\text{Prec}+\text{Recal}}$
Table 7: Reconstruction metric definitions. $P$ and $P^{*}$ are the point clouds sampled from predicted and ground truth mesh.
C.4 Baseline Details
For the results presented in the main paper and supplementary video for Metric3D v2 [26] and Depth Pro [8], we input the ground-truth focal length into their models. The ZoeDepth* [6] model is trained using reproduced code from Depth Anything v1 [75], and we utilize the base model of Depth Anything v1 for conducting experiments. The MSPF results for ARKitScenes dataset are taken from [3], and we retrain it using ScanNet++ [78] training data for testing on Scannet++ with the reproduced code from ARKitScenes.
C.5 Ransac Alignment Details
For monocular depth estimation methods, we perform a post-alignment to ensure fair comparison. We utilize RANSAC alignment to align their output depth with the iPhone LiDAR depth. Specifically, we first resize the output depth to match the dimensions of the iPhone LiDAR depth, then randomly formed several groups of samples. Each group of sample points is used to calculate a scale and shift, followed by voting using all points. The voting threshold is set to the median of the differences between the entire set of numbers and the median(Median Absolute Deviation). Then we apply the scale and shift to the predicted depth to align it with the ground-truth depth. This method is more robust compared to the commonly used polyfit alignment in monocular depth estimation, typically improving the F-score by 8-10% on ScanNet++ dataset.
Figure 16: Qualitative comparison of vehicle LiDAR completion. We include more video results in the supplementary video.
Appendix D Prompting with a vehicle LiDAR
We evaluate our method on the Waymo dataset to assess its performance with vehicle LiDAR. Vehicle LiDAR significantly differs from the LiDAR used in smartphones, as it is generally coarse and consists of X-beam sparse LiDAR (typically 64 beams for Waymo dataset [55]). Therefore, before inputting the data into the network, we perform KNN completion on the vehicle LiDAR depth ($k=4$). We train our model on the Shift dataset [56], a synthetic dataset designed for autonomous driving, which includes RGB and depth data. The LiDAR data is simulated using the approach detailed in the main paper. We evaluate our model on the Waymo dataset. We make comparisons with BPNet [59] in Fig. 16. Our method demonstrates precise depth estimation and we include more video results and street reconstruction results in the supplementary video.
Appendix E Generalized Robotic Grasping Details
Detailed setups.
We control the right arm of a Unitree H1 humanoid robot while fixing its lower body. The task is to grasp the object on the table and put it into the box, one at a time. The object is randomly placed at nearby, middle, and far positions. The robot policy runs at 30 Hz. However, due to overheating issues in our lab environment, the iPhone can only stably capture images at 15 Hz, resulting in the visual input being updated every two control steps.
We first teleoperate the robot to collect 60, 80 trajectories for diffusive objects (red & green cans) and transparent objects (glass bottles); then, we take the diffusive set of data as training set to train ACT [85] policies with different types of visual inputs, including the estimated depth by our model, ARKit depth directly from the iPhone, and also RGB images; during evaluation, we test the grasping performance corresponding to different visual inputs on all objects.
Model architectures.
We use the same network structure with ACT [85] with one image input. ACT policy crops all types of visual input at 480x640 resolution and processes images with a pre-trained ResNet18 backbone[24]. For depth images, the first layer of the pre-trained network is replaced with a 1-channel convolutional network. The pretrained ResNet18 helps enhance the generalization of policy. Without the pretrained parameters, the policy with depth input only grasps the same position.
We include more video results in the supplementary video.

### v3

@@ III-E Implementation Details
We utilize the ViT-large model as our backbone model. The shallow convolutional network comprises two convolutional layers with a kernel size of 3 and a stride of 1, using ReLU as the non-linear activation function. The zero-init project layer is a $1\times 1$ convolution layer.
@@
The irregular range of input depth data can hinder network convergence. To address this, we normalize the LiDAR data using linear scaling to the range [0, 1], based on its minimum and maximum values. The network output is also normalized with the same scaling factor from LiDAR data, ensuring consistent scales and facilitating easier convergence during training. For pattern-agnostic prompting, as the prompt depth points are generally incomplete, the near far range of the prompt depth may not match the near far range of the image depth. In practical, we use the median value of prompt depth points to normalize the prompt depth and make the model output the corresponding normalized depth.
@@
TABLE I: Quantitative comparisons on ScanNet++ dataset. The terms Net., Post. and w/o LiDAR refer to the LiDAR depth usage of models. Methods marked with are finetuned with their released code on ARKitScenes [3] and ScanNet++ [95] datasets.
L1 $\downarrow$ RMSE $\downarrow$ AbsRel $\downarrow$ $\delta_{0.5}$ $\uparrow$ Acc $\downarrow$ Comp $\downarrow$ Prec $\uparrow$ Recall $\uparrow$ F-score $\uparrow$
MSPF 0.0326 0.0975 0.0226 0.9674 0.0772 0.0695 0.6738 0.7761 0.7133
[…]
D.P. 0.0353 0.0983 0.0242 0.9657 0.0820 0.0747 0.6431 0.7234 0.6734
TABLE II: Comparisons on ARKitScenes dataset. The terms Net., Post. and w/o LiDAR refer to the LiDAR depth usage of models, where “Net.” denotes network fusion, “Post.” indicates post-alignment, and “w/o LiDAR” means metric depth.
@@
TABLE III: Quantitative comparisons for depth prompting with completion task on 6 zero-shot testing benchmarks. All methods are re-evaluated on the identical test set and executed on the same machine for a fair comparison of both runtime and performance.
  FPS NYU Scannet DIODE KITTI Waymo ETH3D
AbsRel $\downarrow$ MAE $\downarrow$ AbsRel $\downarrow$ MAE $\downarrow$ AbsRel $\downarrow$ MAE $\downarrow$ AbsRel $\downarrow$ MAE $\downarrow$ AbsRel $\downarrow$ MAE $\downarrow$ AbsRel $\downarrow$ MAE $\downarrow$
  BP-Net 23.81 3.221 0.100 2.678 0.052 3.947 0.251 7.769 1.601 6.203 1.556 5.004 0.478
Marigold-DC 0.05 1.250 0.036 1.190 0.022 0.866 0.030 2.894 0.556 2.875 0.693 1.050 0.066
Omni-DC 7.77 2.164 0.061 2.237 0.042 1.120 0.051 5.670 1.065 7.805 1.864 3.350 0.219
PriorDepthAny. 28.57 1.218 0.036 1.283 0.023 0.471 0.019 2.442 0.573 3.225 0.931 0.791 0.058
Ours 42.02 0.890 0.027 0.900 0.017 0.255 0.011 2.130 0.481 2.900 0.791 1.040 0.054
Figure 7: Qualitative comparisons with SOTA depth completion methods. All the depth maps are normalizing with the same parameters for visualizing.
TABLE IV: Quantitative results on DTU dataset. We evaluate Chamfer Distance of each method on each scene (measured in millimeters) and highlight the best, second-best, and third-best performing metrics in the table.
  COLMAP RealityCapture VolSDF MonoSDF Marigold DepthAny. DepthAny.v2 Metric3D MVSNet IGEV-MVS DUSt3R Murre Ours
  scan24 4.45 4.19 5.24 3.47 4.77 2.99 4.07 5.99 3.16 4.20 2.24 1.31 0.82
scan37 4.67 3.85 5.09 3.61 7.61 3.77 4.33 5.26 4.36 5.05 3.43 3.14 2.52
scan40 2.51 2.26 3.99 2.10 5.34 3.01 2.54 4.13 2.59 3.47 2.77 1.53 1.00
scan55 1.90 2.49 1.42 1.05 4.75 2.10 4.07 4.22 2.35 2.79 2.40 0.92 0.61
scan63 2.81 3.49 5.10 2.37 4.07 2.98 3.05 3.14 1.88 3.05 3.71 1.19 1.11
scan65 2.92 3.97 4.33 1.38 4.15 4.10 3.61 5.00 3.28 3.16 3.22 1.98 1.16
scan69 2.12 1.91 5.36 1.41 7.38 2.73 3.16 4.20 2.32 3.16 2.49 1.25 0.84
scan83 2.05 2.49 3.15 1.85 6.00 2.38 3.29 4.78 1.88 3.17 3.34 2.16 1.39
scan97 2.93 2.37 5.78 1.74 5.38 4.17 4.86 4.54 1.87 3.62 2.71 1.40 1.19
scan105 2.05 2.27 2.07 1.10 5.26 3.45 2.79 4.85 2.06 4.16 2.84 1.07 0.75
scan106 2.01 2.90 2.79 1.46 5.36 4.40 5.24 5.44 2.25 3.09 2.69 1.26 0.91
scan110 N/A 4.60 5.73 2.28 5.51 1.97 5.29 4.69 1.92 5.97 3.08 0.95 1.01
scan114 1.10 1.38 1.20 1.25 4.65 2.53 2.41 7.82 1.40 2.49 2.20 0.82 0.47
scan118 2.72 2.57 5.64 1.44 5.63 2.94 3.01 4.36 2.43 3.23 2.52 1.18 0.70
scan122 1.64 1.76 6.20 1.45 5.99 2.88 3.57 6.78 1.94 1.90 2.51 1.07 0.78
Average 2.56 2.84 4.21 1.86 5.46 3.09 3.69 5.01 2.38 3.50 2.81 1.42 1.02
Figure 8: Qualitative results on DTU dataset. All the methods use the same 3 input views.
TABLE V: Quantitative results on Replica dataset. We evaluate F-Score of each method on each scene and highlight the best, second-best, and third-best performing metrics in the table.
  room_1 office_0 office_2 Average
  MonoSDF - - - 0.86
Marigold 0.61 0.52 0.58 0.57
SparseDC 0.25 0.29 0.23 0.25
MVSNet 0.54 0.73 0.56 0.61
IGEV-MVS 0.84 0.85 0.78 0.82
Murre 0.84 0.90 0.82 0.85
Ours 0.91 0.92 0.86 0.90
IV Experiments
IV-A Experimental Setup
Prompting a depth foundation model with depth priors can be categorized into two types of tasks: depth upsampling and depth completion. For depth prompting with upsampling task, we mainly conduct experiments on the HyperSim synthetic dataset [60] and two real-world datasets: ScanNet++ [95] and ARKitScenes [3], which provide iPhone RGB-LiDAR data ($192\times 256$ resolution) and annotated depth from a high-power LiDAR ($1440\times 1920$ resolution). We follow the suggested training and evaluation protocol in [3] for ARKitScenes, where 40K images are used for training and 5K images for evaluation. For the ScanNet++ dataset, we randomly select 20 scenes from its 50 validation scenes, amounting to approximately 5K images for our validation and the training set are from its 230 training scenes, containing about 60K images. To ensure a fair comparison, we additionally train a model with HyperSim training set to achieve zero-shot testing on ScanNet++ and ARKitScenes datasets.
For depth prompting with completion task, we train our model using common synthetic monocular depth datasets including HyperSim [60], VKITTI2 [10], TartanAir [16], UrbanSyn [23], DynamicReplica [35], datasets. We evalute our method on a diverse set of six zero-shot datasets: three indoor scene datasets, NYUv2 [65], ScanNet [16], and DIODE [72]; one outdoor scene dataset, ETH3D [64]; and two autonomous driving datasets, KITTI [22] and Waymo [67]. All the datasets are zero-shot for our method. Prompt depth maps, serving as input, were generated by randomly selecting 1,500 points from the sensor-acquired depth maps of each dataset. We report depth accuracy metrics including AbsRel and L1 error, which stand for relative percentage error and mean absolute error in distance, respectively.
IV-B Comparisons with the State of the Art
@@
For the depth completion task, we compare our method with four state-of-the-art depth completion approaches [70, 78, 103, 74] across six widely used benchmarks [65, 16, 72, 64, 67, 22]. [70, 78, 103] use network fusion and [74] proposes a test time optimization. These datasets are captured with different types of sensors and exhibit extremely diverse patterns, as shown in Fig.˜7. Tabs.˜III and 7 demonstrates that our method outperforms previous state-of-the-art methods in both accuracy and inference speed, making it a highly promising solution for practical applications. Note that our method even presents better results than Marigold-DC [74] which requires a long optimization process for each inference. All the benchmarks are zero-shot for our method, demonstrating the strong generalization ability of our approach across diverse scenarios and sensor patterns.
3D reconstruction is an important downstream task for depth estimation. A recent trend is to use SfM depth as a condition for monocular depth models [25], combining the 3D consistency of SfM points with the strong prior of monocular depth estimation. Our pattern-agnostic prompting is highly suitable for this setting. To further demonstrate the effectiveness of our pattern-agnostic prompting, we follow the experimental setup of Murre [25], using SfM depth to prompt our model. Specifically, we evaluate our model on the DTU [34] and Replica [66] datasets. We use Detector-free SfM [29] to obtain the SfM depth for both datasets. These sparse depth maps are fed into our method to obtain dense depth predictions. We then run TSDF to obtain the final mesh results. All the evaluation settings are aligned with Murre [25] to make faire comparisons. As shown in Tabs.˜IV, V and 8, our method presents clear improvements over baselines. Note that our method is nerver finetuned on Structure-from-Motion datasets or DTU MVS datasets [34], while previous methods [25, 86] train their model on SfM or DTU MVS datasets, respectively. These comparisons highlights the effectiveness of our pattern-agnostic prompting, providing a new scalable solution for 3D reconstruction field which only needs monocular depth training.
TABLE VI: Quantitative ablations on ARKitScenes and ScanNet++ datasets. Please refer to Sec.˜IV-C for detailed descriptions.
@@
(k) Depth Pro as foundation 0.0169 0.0150 0.0754 0.0676 0.7202
(l) Depth Pro 0.1225 0.1038 0.0904 0.0760 0.6187
TABLE VII: Quantitative ablations for prompting strategies. The upsampling prompt input is $30\times 40$ low-res depth and completion prompt input is 12001200 random points.
  Upsampling Setting Completion Setting
AbsRel $\downarrow$ MAE $\downarrow$ AbsRel $\downarrow$ MAE $\downarrow$
  Upsampling Prompting Strategy 0.89 0.109 1.67 0.151
Pattern-agnostic Prompting Strategy 1.09 0.125 1.09 0.124
IV-C Ablations and Analysis
@@
We ablate the importance of prompting a depth foundation model with two experiments: 1) Removing the prompting. Tab.˜VI(b) shows a significant performance drop. 2) Removing the foundation model initialization [93]. Tab.˜VI(c) shows a noticeable performance decline.
@@
We study different designs: AdaLN, Cross-attention, and ControlNet as discussed in Sec.˜III-B. Tab.˜VI(d,e,f) reveals that ControlNet performs best but still falls short of our method.
@@
We initially incorporate ARKitScenes data, which only enhances performance on ARKitScenes (Tab.˜VI(g)). Then we add ScanNet++, which improves results on both ARKitScenes and ScanNet++ (Tab.˜VI(h)). However, the depth visualization remains less than ideal (Fig.˜3). Tab.˜VI(i) show that direct supervision with pseudo GT depth from reconstruction methods decreases performance. Ultimately, employing the edge-aware depth loss that utilizes pseudo GT depth and FARO annotated GT achieves comparable performance with  Tab.˜VI(h) but with superior thin structure depth performance as shown in Fig.˜3.
Prompting strategies.
In this work, we propose two prompting strategies tailored for depth upsampling and completion tasks, respectively. To assess their effectiveness, we conduct experiments on the HyperSim test set, evaluating both models under upsampling and completion scenarios. As shown in Tab.˜VII, each prompting strategy performs best in its designated setting. The pattern-agnostic prompting achieves comparable results in both tasks, while the upsampling prompting excels specifically in upsampling. These results indicate that both strategies are necessary: the upsampling prompting is optimal for upsampling, whereas the pattern-agnostic prompting offers broader applicability across different tasks.
Figure 9: Outdoor reconstruction by taking the vehicle LiDAR as prompt.
Replacing the foundation model.
Since our model is a general design for DPT, it can be easily adapted to other depth foundation models that also utilize the DPT structure, such as Depth Pro [8]. Our experiments demonstrate that it significantly enhances the performance of Depth Pro, as shown in Tabs.˜II and VI-(k), although it does not outperform our choice of Depth Anything.
Figure 10: Zero-shot testing on diverse scenes.
@@
Our upsampling model with ViT-L runs at 20.4 FPS for an image resolution of $768\times 1024$ on a A100 GPU. As ARKit6 supports 4K image recording, we test our model at a resolution of $2160\times 3840$ and achieve 2.0 FPS. Note that our model can also be implemented with ViT-S, where the corresponding speeds are 80.0 and 10.3 FPS. More testing results can be found in the supp. Our completion model with ViT-L achieves a runtime of 18.2 FPS at a resolution of $768\times 1024$. This is slightly slower than the upsampling model, as the prompting attention block introduces additional computational overhead, accounting for approximately 12% of the total computation of the depth foundation model.
IV-D Zero-shot Testing on Diverse Scenes
Although our model is trained on indoor scenes, it generalizes well to various scenarios, including new rooms, gyms with thin structures, poorly lit museums, human and outdoor environments, as shown in Fig.˜10, highlighting the effectiveness of prompting a depth foundation model.
IV-E Application: Street Reconstruction
Our consistent and scale-accurate depth estimation benefits the indoor, object-level 3D reconstruction as shown in Tabs.˜I, 5 and 8. Besides, the prompt of our model can be easily replaced with vehicle LiDAR, which enables our model to achieve large-scale outdoor scene reconstruction as shown in Fig.˜9. We detail the setup in the supp.
IV-F Application: Generalized Robotic Grasping
@@
We set up a robotic platform to test our model in generalized robotic manipulation (Fig. 11), which typically requires depth or RGB as observations. Good depth estimation enhances the generalization ability because it accurately describes the 3D information of surroundings [99, 32]. Specifically, we train an ACT policy [102] to grasp various objects into the box, using different types of input signals such as RGB, LiDAR, and depth data from our model. We empirically find that our model generalizes well to unseen objects like transparent and specular objects when trained on diffusive objects, outperforming RGB and LiDAR inputs as shown in Tab.˜VIII. This is because RGB is dominated by color, which leads to poor generalization across objects, and the iPhone LiDAR depth is noisy and lacks the capability to perceive transparent objects. Please refer to the supp. for detailed setup descriptions and videos.
TABLE VIII: Grasping success rate on various objects. Three numbers indicate objects placed at near, middle, and far positions. The grasping policy is trained on diffusive and tested on all objects.
@@
V Conclusion and Discussions